If it takes 3 cups of sugar to make cookies for 40 children, how many cups of sugar will it take to make cookies for 600 students?
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

3/40=x/600
40x=1800
x=1800/40
x=45 cups

Answer by stanbon(75887)

If it takes 3 cups of sugar to make cookies for 40 children, how many cups of sugar will it take to make cookies for 600 students?
----
cups/children = cups/children
----
x/600 = 3/40
-----
x = 600(3/40)
----
x = 45 cups
------------
Cheers,
Stan H.
----------